OpenStudy (anonymous):

How do I figure out the area and perimeter of a square that is 17 yd wide and 10 yd long

OpenStudy (kkutie7):

perimeter= l+w+l+w =2l+2w

OpenStudy (just_one_last_goodbye):

then it's not a square ._.

OpenStudy (just_one_last_goodbye):

a square are all the same lengths

OpenStudy (kkutie7):

Area=l*w

OpenStudy (anonymous):

Well the formula for area is A=lw With means you multiple A=17 x 10 And so Area is 170 yards and perimiter formula is P=l +w so perimeter is 27 yards

OpenStudy (just_one_last_goodbye):

A=170

OpenStudy (anonymous):

rectangle lol, thank you

OpenStudy (anonymous):

P=l+w+l+w so 27 + 27= 54 yards

OpenStudy (anonymous):

or just 27 x 2

OpenStudy (just_one_last_goodbye):

I got the answer by multiply 17 * 10 = 170

OpenStudy (anonymous):

yep l x w =A